OpenStudy (anonymous):

The formula for degrees Celsius is: C =5/9 (F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is

OpenStudy (anonymous):

C*(9/5) = F-32 (9/5)C + 32 = F